A club sold fruitcakes for a fund raising project. The two-pound cakes sold for $5 each and the five-pound cakes sold for $11 each. The club sold 323 pounds of cakes and received $749. How many of each did they sell?
Answer by stanbon(75887)

A club sold fruitcakes for a fund raising project. The two-pound cakes sold for $5 each and the five-pound cakes sold for $11 each. The club sold 323 pounds of cakes and received $749. How many of each did they sell?
-------------------
Quantity Eq.::: 2t + 5f = 323 lbs
---
Value Eq: 5(t) + 11(f) = 749 dollars
----------------------------------------
Multiply thru 1st by 5
Multiply thru 2nd by 2
--------------------------
10t + 25f = 5*323
10t + 22f = 2*749
----------------------
Subtract 2nd from 1st and solve for "f":
3f = 117
f = 29 (# of 5-pound cakes)
---
Use 2t + 5f = 323 to solve for "t":
2t + 5*29 = 323
2t = 178
t = 89 (# of 2-pound cakes)
==============================
Cheers,
Stan H.